OpenStudy (kenshin):

Hi all, I have encountered a contradiction which I cannot resolve by myself, so someone please help me out here: we know that the formula for the inverse of a 2x2 A if A is [a b; c d] then the inverse of A is (1/(ad-bc)) * [d -b; -c a] but from my textbook it says: det (A^-1) should = 1/(det(A)); because of: det(A*A^-1) = det(I) = 1 = det(A) * det(A^-1) and that makes det(A^-1) = 1/det(A) now if i was to use the 2x2 matrix as an example and do det(A), i get ad-bc, and det(A^-1) = det((1/(ad-bc)*[d -b; -c a]) is just 1! since you get (da-bc)/(ad-bc). That sure isn't right?

OpenStudy (amistre64):

ab cd inverse, isnt the inverse divided by the det?

OpenStudy (dan815):

it has to be 1

OpenStudy (amistre64):

hard to parse your postings ... but i see where you have it

OpenStudy (mathmate):

remember det(kA)=k^2 det(A)

OpenStudy (kenshin):

the inverse from the 2x2 inverse formula makes it (1/(ad-bc)) * [d -b] [-c a] taking the det of the above inverse makes it 1. but det (A) is just ad-bc, so according to the book, 1/(ad-bc) should be the det of the above inverse...

OpenStudy (dan815):

no thats not how it works

OpenStudy (dan815):

u cannot multiply scalar like that look what mathmate said

OpenStudy (dan815):

think about exapanding the scalar into your matrix and then what happens when u take determinant

OpenStudy (kenshin):

so dan Det(A^-1) does not equal 1/(det(A))?

OpenStudy (dan815):

it does

OpenStudy (amistre64):

what is n/n^2 ?

OpenStudy (mathmate):

*determinant(kA)=k^n determinant(A) In this case, n=2, so determinant((1/(ad-bc))*A) =(1/(ad-bc))^2 determinant (A) =(ad-bc)/(ad-bc)^2 =1/(ad-bc)

OpenStudy (amistre64):

let det(A) = n nd -nb -nc na whats our determinant of the inverse?

OpenStudy (dan815):

http://prntscr.com/8zjj4h

OpenStudy (kenshin):

so i can't take the 1/(ad-bc) out of the bracket as a scalar for the det of the inverse? i remember det(kA) = kdet(A) though...

OpenStudy (dan815):

no thats what im trying to tell u

OpenStudy (dan815):

look at amistres post.. when u expand it over a 2by 2 matrix now u multiply the constant twice

OpenStudy (amistre64):

slight mistype ... let 1/det(A) = n :)

OpenStudy (dan815):